# On the Capacity of Self-Attention

### v1

@@ 7 Lower Bound on Relational Graph Recognition
A parameter choice $\big(\{W_{Q}^{(k)},W_{K}^{(k)}\}_{k=1}^{h},\tau\big)$ recovers a graph $G=(V,E)$ with margin $\gamma>0$ if, for every ordered pair $(u,v)$ with $u\neq v$,
$(u,v)\in E\Rightarrow\mathbf{x}_{u}^{\top}M\,\mathbf{x}_{v}\geq\tau+\gamma,\qquad(u,v)\notin E\Rightarrow\mathbf{x}_{u}^{\top}M\,\mathbf{x}_{v}\leq\tau-\gamma,$
where $M=\sum_{k}W_{Q}^{(k)}W_{K}^{(k)\top}$.
Because the decision rule is homogeneous in $(M,\tau)$, we fix scale by requiring
$M=UV^{\top},\quad\|U\|_{F}\leq 1,\ \|V\|_{F}\leq 1,\ \ |\tau|\leq 1.$ (6)
Under (6) and $\|\mathbf{x}_{u}\|_{2},\|\mathbf{x}_{v}\|_{2}\leq 1$, the score map is $1$-Lipschitz in Frobenius norm:
$\big|\mathbf{x}_{u}^{\top}UV^{\top}\mathbf{x}_{v}-\mathbf{x}_{u}^{\top}\tilde{U}\tilde{V}^{\top}\mathbf{x}_{v}\big|\ \leq\ \|U-\tilde{U}\|_{F}+\|V-\tilde{V}\|_{F},\qquad\text{and }\ |(\tau-\tilde{\tau})|\ \text{adds linearly.}$ (7)
Therefore any perturbation of $(U,V,\tau)$ of radius at most $\gamma/4$ preserves all pairwise signs, and hence the entire edge set, on length‑2 contexts.
@@
Fix $m\in\mathbb{N}$ and $m^{\prime}\in\{0,\dots,m(m\!-\!1)\}$. Suppose an attention mechanism of the form (3)–(4), with item embeddings $\|\mathbf{x}_{v}\|_{2}\leq 1$, can recover every graph in $\mathcal{G}_{m,m^{\prime}}$ with margin $\gamma\in(0,1)$. Then there exists a constant $c(\gamma)>0$ such that
$d_{\text{model}}\;D_{K}\ \ \geq\ \ c(\gamma)\,\log\binom{m(m-1)}{m^{\prime}}\ -\ O(1).$ (8)
Equivalently,
$D_{K}\ \ =\ \ \Omega\!\Bigg(\frac{\log\binom{m(m-1)}{m^{\prime}}}{d_{\text{model}}}\Bigg).$ (9)
@@
Under the normalization (6) and Lipschitz property (7), the parameter set $\mathbb{B}=\{(U,V,\tau):\|U\|_{F},\|V\|_{F},|\tau|\leq 1\}$ admits an $\varepsilon$-net of radius $\varepsilon=\gamma/4$ of size at most
@@
for absolute constants $C,C^{\prime}>0$. Each net point induces a unique labeling of the N=m(m1)N=m(m-1) ordered pairs by the margin, hence at most $N_{\mathrm{cov}}(\varepsilon)$ distinct edge sets can be realized. Since the mechanism must realize all $\binom{N}{m^{\prime}}$ edge sets of size $m^{\prime}$, we obtain $\binom{N}{m^{\prime}}\leq N_{\mathrm{cov}}(\varepsilon)$, which rearranges to (8)–(9). ∎
Equivalent finite‑precision statement.
If each real parameter in $\{W_{Q}^{(k)},W_{K}^{(k)},\tau\}$ has $b=\Theta(1)$ effective bits after normalization (e.g., due to quantization or stochastic rounding), then the parameter budget contains at most $B=b\,(2\,d_{\text{model}}D_{K}+1)$ bits and thus can realize at most $2^{B}$ distinct edge sets. Requiring $2^{B}\geq\binom{N}{m^{\prime}}$ gives the same conclusion as (9) with an explicit constant $1/(2b)$. Under the margin model above, one may take $b=\Theta(\log(1/\gamma))$.
Bounds for specific cases.
This demonstrates the following results:
  • Exactly $m^{\prime}=m$ edges (such as permutation graphs): $D_{K}=\Omega\Big(\frac{m^{\prime}\,\log m}{d_{\text{model}}}\Big).$
  • Dense regime with $m^{\prime}=\Theta(m^{2})$: $D_{K}=\Omega\Big(\frac{m^{\prime}}{d_{\text{model}}}\Big)$.
  • Sparse regime with $m^{\prime}=O(m^{2-\epsilon})$ for some positive constant $\epsilon$: $D_{K}=\Omega\Big(\frac{m^{\prime}\,\log m}{d_{\text{model}}}\Big)$.
@@ 8 Limitations
Theory. Our constructive upper bounds are tight only under a mild degree–skew assumption (e.g., $\Delta/d_{\text{avg}}\leq m/d_{\text{model}}$). When this condition is violated—or in very dense graphs—the present upper bounds do not match the information-theoretic lower bound. Moreover, our sufficiency results rely on geometric properties of the embeddings (near-orthogonality / restricted self-incoherence). While we verify these conditions for idealized random embeddings, we do not quantify how embeddings learned in trained models satisfy (or deviate from) these properties, nor the effect of such deviations on the constants in our bounds.
Experiments. Empirically we restrict to permutation graphs (out-degree =1=1) and do not test graphs with larger out-degree. We also evaluate a QK-only layer and thus do not experimentally validate modeling assumptions such as adding a value pathway or incorporating a softmax decision rule. Our search over $D_{K}$ is coarse because $D_{K}$ is tuned in multiples of the head count; there remains room to probe all cross-points in the scanned range and to explore substantially larger dimensions, but we did not do so due to compute limits. Finally, while capacity appeared largely insensitive to context length within $\ell\in\{16,32\}$, we did not study much longer contexts for the same reason.
@@ A.1 Experimental implementation Details
$z_{pq}\;=\;\alpha\big(S_{\max}(p,q)-\tau\big),\qquad\mathcal{L}\;=\;\tfrac{1}{|\mathcal{C}|^{2}}\!\sum_{p,q}\!\Big[\underbrace{\mathrm{softplus}(-z_{pq})y_{pq}}_{\text{positive term}}\cdot\underbrace{\mathrm{pos_{weight}}}_{=\ell-1}+\mathrm{softplus}(z_{pq})(1-y_{pq})\Big],$
@@ A.2 Result details
A.3 Sensitivity to context length.
@@
[…]
@@ Appendix B Further Details on Our Explicit Constructions
We here provide additional details on our explicit constructions from Section 4. We start with the easiest case - permutation graphs with no embedding. This result is subsummed by the second construction, so is only included as a warm up for the more general case.
@@ B.1 Construction I: Permutation Graphs with One‑Hot Embeddings
3:
4:Construct Key Matrix: Draw $W_{K}\in\mathbb{R}^{m\times d_{k}}$ with i.i.d. entries $(W_{K})_{jl}\sim\text{Bernoulli}(p)$.
5:   For each node $v_{j}$, the key is $\mathbf{k}_{j}=\mathbf{e}_{j}W_{K}$.
6:
7:Construct Query Matrix: For each node $v_{i}$, set its query $\mathbf{q}_{i}=\mathbf{k}_{\pi(i)}$.
8:   This is equivalent to setting the $i$-th row of $W_{Q}$ to be the $\pi(i)$-th row of $W_{K}$.
9:
10:Set Threshold: $\tau=\frac{p+p^{2}}{2}\,d_{k}$.
@@
$\Pr\!\big[S_{i,\pi(i)}\leq\tau\big]\;\leq\;\exp\!\Big(-\tfrac{\mu_{1}\delta_{1}^{2}}{2}\Big)\quad\text{where}\quad\delta_{1}=1-\tfrac{\tau}{\mu_{1}}=\tfrac{1-p}{2},$
@@
$\Pr\!\big[S_{ij}\geq\tau\big]\;\leq\;\exp\!\Big(-\tfrac{\mu_{2}\delta_{2}^{2}}{2+\delta_{2}}\Big)\quad\text{where}\quad\delta_{2}=\tfrac{\tau}{\mu_{2}}-1=\tfrac{1-p}{2p},$
@@
Lemma 4.1 immediately now yields correctness for every context, independent of context length. Our lower bound from Section 7 for this case is $\Omega(\frac{m^{\prime}}{d_{\text{model}}}\log m)=\Omega(\log m)$. Our construction achieves an upper bound of $d_{k}=O(\log m)$, demonstrating that the bound is tight for this class of problems. Also note that the threshold proof is identical if softmax is used; see Appendix C.

### v3

@@ 9 Lower Bounds on Relational Graph Recognition
A parameter choice $\big(\{U_{k},V_{k}\}_{k=1}^{h},\tau\big)$ recovers a graph $G=(V,E)$ with margin $\gamma>0$ if, for every ordered pair $(u,v)$ with $u\neq v$,
$(u,v)\in E\Rightarrow S(u,v)\geq\tau+\gamma,$
$(u,v)\notin E\Rightarrow S(u,v)\leq\tau-\gamma,$
with $S$ given by (26).
@@
Fix $m\in\mathbb{N}$ and $m^{\prime}\in\{0,\dots,m(m\!-\!1)\}$. Suppose a single self‑attention QK channel with total key dimension $D_{K}=\sum_{k=1}^{h}d_{k}$ and embeddings $\mathbf{x}_{v},v\in V$ can recover every graph in $\mathcal{G}_{m,m^{\prime}}$ with margin $\gamma\in(0,1)$ on contexts of any length $\ell\geq 2$. Further, assume that the recovery is achievable with parameters and embeddings whose norms are bounded by a universal constant independent of $m$. Then there exists a constant $c(\gamma)>0$ such that
$d_{\text{model}}\,D_{K}\ \ \geq\ \ c(\gamma)\,\log\binom{m(m-1)}{m^{\prime}}\ -\ O(1),$ (24)
or equivalently
$D_{K}\ \ =\ \ \Omega\!\Bigg(\frac{\log\binom{m(m-1)}{m^{\prime}}}{d_{\text{model}}}\Bigg)\ =\ \Omega\!\Big(\tfrac{m^{\prime}}{d_{\text{model}}}\log\tfrac{m^{2}}{m^{\prime}}\Big).$ (25)
@@
We first point out that we can focus only on length‑2 contexts. Uniform correctness for RGR requires that, for every ordered pair $(u,v)\in V\times V$, the decision $(u,v)\in E$?” is the same in every context containing $u$ and $v$. In particular it must be correct in the length‑2 context $\mathcal{C}=(u,v)$. Hence any lower bound proved using only length‑2 contexts applies to the full problem.
For head $k$, write the per‑pair score on a length‑2 context as
$s_{k}(u,v)\;=\;\mathbf{x}_{u}^{\top}A_{k}\,\mathbf{x}_{v},$
$A_{k}\;=\;W_{Q}^{(k)}W_{K}^{(k)\top},\quad\mathrm{rank}(A_{k})\leq d_{k}.$
Define the final score $S(u,v)=\max_{k}s_{k}(u,v)$, and decide
$(u,v)\in E\;\Leftrightarrow\;S(u,v)>\tau.$ (26)
We factor each $A_{k}=U_{k}V_{k}^{\top}$ with $U_{k},V_{k}\in\mathbb{R}^{d_{\text{model}}\times d_{k}}$, and concatenate $U=[U_{1}|\cdots|U_{h}]\in\mathbb{R}^{d_{\text{model}}\times D_{K}}$, $V=[V_{1}|\cdots|V_{h}]\in\mathbb{R}^{d_{\text{model}}\times D_{K}}$. The following normalization assumption is without loss of generality by scale homogeneity of the decision rule (simultaneously scaling $(U,V)$ and $\tau$), and our assumption of bounded norms of our embeddings and parameters:
$\begin{split}&\|U\|_{F}\leq 1,\quad\|V\|_{F}\leq 1,\\ &|\tau|\leq 1,\quad\|\mathbf{x}_{v}\|_{2}\leq 1\quad\forall v\in V.\end{split}$ (27)
Note that without the bounded norms assumption, this normalization step could introduce a dependence on $m$ into the margin $\gamma$, which in turn would carry into the final lower bound.
Lemma 9.3 (Lipschitz property of the Decision Function).
[…]
Sketch.
For each head, using $\|\mathbf{x}_{u}\|_{2},\|\mathbf{x}_{v}\|_{2}\leq 1$ and $\|U_{k}\|_{F},\|V_{k}\|_{F}\leq 1$ (implied by (27)),
$\big|s_{k}-\tilde{s}_{k}\big|$ $=\big|\mathbf{x}_{u}^{\top}(U_{k}V_{k}^{\top}-\tilde{U}_{k}\tilde{V}_{k}^{\top})\mathbf{x}_{v}\big|$
$\leq\|U_{k}-\tilde{U}_{k}\|_{F}+\|V_{k}-\tilde{V}_{k}\|_{F}.$
[…]
We also note that the same Lemma holds (using a similar proof) if we instead use either a log-sum-exp aggregation rule or a summation aggregation rule: our lower bounds hold for those cases as well.
We also note that the same covering-number argument applies to other common aggregations (including $\sum_{k}$ and log-sum-exp across heads), and specifically to our more standard softmax model in which each head applies a softmax over the context and we sum the resulting attention weights across heads before thresholding; see the paragraph “Softmax attention” below.
From our normalization assumption (27), every parameterization lies in the compact ball
$\mathbb{B}\ :=\ \big\{(U,V,\tau):\ \|U\|_{F}\leq 1,\ \|V\|_{F}\leq 1,\ |\tau|\leq 1\big\},$
whose radius is at most 1. Therefore, under the normalization (27) and Lemma 9.3, the parameter set admits an $\varepsilon$-net (in Frobenius metric for $U,V$ and absolute value for $\tau$) of radius $\varepsilon=\gamma/4$ and size at most
@@
for absolute constants $C,C^{\prime}>0$. Each net point induces a unique labeling of the N=m(m1)N=m(m-1) ordered pairs by the $\gamma$‑margin, because any perturbation of radius $\varepsilon$ preserves all signs by (28). Therefore at most $N_{\mathrm{cov}}(\varepsilon)$ distinct edge sets can be realized at margin $\gamma$. Since the mechanism must realize all $\binom{N}{m^{\prime}}$ edge sets of size $m^{\prime}$, we obtain $\binom{N}{m^{\prime}}\leq N_{\mathrm{cov}}(\varepsilon)$, which rearranges to (24)–(25). ∎
Softmax attention.
Theorem 9.2 also applies to the softmax version of our model. Uniform recovery requires the decision for a fixed ordered pair $(i,j)$ to be identical in every context containing $i$ and $j$, so it suffices to analyze the length-$2$ context $\mathcal{C}=(i,j)$. For $\ell=2$ each head reduces to a two-way softmax:
$p_{ij}^{(k)}$ $=\frac{\exp(L_{ij}^{(k)})}{\exp(L_{ii}^{(k)})+\exp(L_{ij}^{(k)})}$ (29)
$=\sigma\!\big(L_{ij}^{(k)}-L_{ii}^{(k)}\big),\qquad\|\sigma^{\prime}\|_{\infty}\leq\tfrac{1}{4},$
[…]
$|A_{ij}-\tilde{A}_{ij}|\;\leq\;\tfrac{1}{2}\big(\|U-\tilde{U}\|_{F}+\|V-\tilde{V}\|_{F}\big).$ (30)
Thus the decision map remains $O(1)$-Lipschitz, so the same $\varepsilon$-net/margin-stability argument as in the proof of Theorem 9.2 yields the identical lower bound (25).
Bounds for specific regimes.
The theorem implies:
  • Exactly $m^{\prime}=m$ edges (e.g., permutation graphs): $D_{K}=\Omega\!\big(\tfrac{m^{\prime}\,\log m}{d_{\text{model}}}\big)$.
  • Dense regime with $m^{\prime}=\Theta(m^{2})$: $D_{K}=\Omega\!\big(\tfrac{m^{\prime}}{d_{\text{model}}}\big)$.
  • Sparse regime with $m^{\prime}=O(m^{2-\epsilon})$ for some $\epsilon>0$: $D_{K}=\Omega\!\big(\tfrac{m^{\prime}\,\log m}{d_{\text{model}}}\big)$.
@@ 10 Limitations
Our theoretical guarantees on $D_{K}$ are not uniformly tight: for max-over-heads we asymptotically match the information-theoretic lower bound for permutation graphs and broad sparse regimes, whereas gaps remain for very dense and highly degree-imbalanced graphs. For softmax our sufficient condition incurs an extra log factor whose necessity is unclear. Since a central motivation is to show the multi-head advantage does not rely on “attending to multiple targets at once,” permutation graphs are the most diagnostic setting, and this is where our bounds and evidence are strongest. Our analysis assumes a controlled embedding geometry to quantify interference; we partially bridge this by validating the phenomenon in a full single-layer Transformer block with frozen GPT-2 embeddings, and we expect realistic embeddings to create more (not less) interference via concentration. While learned embeddings could improve constants, our lower bounds (which still apply with learned embeddings) together with our near-matching upper bounds limit how much they can help. Empirically, we focus on controlled single-layer (largely synthetic) tasks, so we do not model multi-layer iterative routing, or end-to-end training dynamics.
@@ A.1 Experimental implementation details
$z_{pq}\;=\;\alpha\big(S_{\max}(p,q)-\tau\big),$
$\begin{split}\mathcal{L}\;=\;\tfrac{1}{|\mathcal{C}|^{2}}\!\sum_{p,q}\!\Big[&\underbrace{\mathrm{softplus}(-z_{pq})y_{pq}}_{\text{positive term}}\cdot\underbrace{\mathrm{pos_{weight}}}_{=\ell-1}\\ &+\;\mathrm{softplus}(z_{pq})(1-y_{pq})\Big],\end{split}$
@@ A.2 Result details
A.3 Experiments for Denser graphs
We next consider denser graphs in the same max-over-heads model variant. Our theoretical results demonstrate tight or nearly tight asymptotic bounds for a broad class of denser graphs; this section seeks to validate and refine those results through experiments in a setting that mirrors our theoretical model. In particular, we examine $r$ regular graphs, and study the impact of scaling $r$.
Unless noted, the architecture, training loop, optimizer, early‑stopping protocol, evaluation metric, and data split are identical to the permutation‑graph experiments. The only substantive differences are:
  • Task / graph family. We replace the permutation graph (out‑degree =1=1) with a directed rr‑regular graph on $m$ nodes (every node has exactly $r$ out‑neighbors and $r$ in‑neighbors; no self‑loops, no multi‑edges). Concretely, the graph is sampled as the union of $r$ random perfect matchings (“layers”), each built by a randomized greedy permutation under “forbidden” constraints to avoid self‑loops and duplicates; we verify that all in‑degrees equal $r$.
  • Labeling. For a context $X_{\mathcal{C}}$, an ordered pair $(p\!\to\!q)$ is positive iff $q\in N^{+}(p)$ and both $p,q\in\mathcal{C}$. Thus each row can contain up to $r$ positives (vs. at most one in the permutation case). Prediction still uses the per‑head scores $S^{(k)}$, an elementwise max across heads $S_{\max}$, and a single learned global threshold $\tau$.
  • Context construction (fixed target‑in‑context rate). We retain the same target‑in‑context rate $\rho$, but now implement it over sources: we draw $b\!\sim\!\mathrm{Binom}(\ell,\rho)$ sources from the sampled context and ensure that for each selected source $u$ at least one out‑neighbor of $u$ is included in the context. When r=1r{=}1 this reduces to the permutation procedure.
  • Class‑imbalance weighting. Because the number of positives per context grows with $r$, we replace the fixed positive weight $(\ell{-}1)$ used for permutations with a batch‑adaptive weight
    $w_{+}\;=\;\frac{\#\text{negatives}}{\#\text{positives}}$ (31)
    computed per context inside the same logistic loss on logits $\alpha\!\left(S_{\max}-\tau\right)$ (with the same $\alpha$).
Everything else (frozen, normalized Gaussian node embeddings; idealized attention with no softmax/value path; elementwise max across heads; single learned $\tau$; AdamW with the same hyperparameters; validation/test protocol; and the micro‑F1 reporting criterion) is as in the baseline/core permutation‑graph experiments. We also here test the specific case of $m=128$ and $d_{\text{model}}=32$ and scale $r$ from 1 to 32. Given the consistency of our results across different values of $m$ and $d_{\text{model}}$, we believe that this serves as a good proxy for more general behavior.
Findings and impact (denser graphs).
Moving from permutation graphs (r=1r{=}1) to denser, $r$-regular graphs preserves the qualitative behavior of the capacity transition and sharpens several quantitative predictions about how the key–query budget should scale.
Sharp capacity transition persists with density.
@@ A.3 Experiments for Denser graphs
Capacity is governed by edges, not vertices.
Normalizing by the edge budget
$x\;=\;\frac{m^{\prime}}{d_{\text{model}}}\log_{2}m\;=\;\frac{r\,m}{d_{\text{model}}}\log_{2}m,$ (32)
the empirically minimal $D_{K}$ required for high accuracy collapses to a single linear trend across degrees $r\in\{1,2,4,8,16\}$ (Figure 12). A single proportionality constant fits all densities:
$D_{K}^{\star}\;\approx\;0.46\,x,$ (33)
so, as in our theoretical results, the total key dimension tracks $m^{\prime}$ much more tightly than $m$. Intuitively, the model’s where-to-attend budget must scale with the number of encoded relationships; increasing vertices without adding edges exerts far less pressure on $D_{K}$.
Head count scales with edge density.
The number of heads at the empirical threshold grows with graph density and is well predicted (up to a small constant factor) by the edge-normalized ratio $m^{\prime}/d_{\text{model}}$ (Figure 13). In our runs with m=128m{=}128 and $d_{\text{model}}{=}32$, the head count that achieves the smallest passing $D_{K}$ increases roughly linearly with $r$ and stays close (within a factor of $\sim\!2$) to the theoretical target $h^{\star}\propto m^{\prime}/d_{\text{model}}=(r\,m)/d_{\text{model}}$. This reinforces a capacity-based rationale for multi-head attention: as more edges are superposed in the compressed embedding space, distributing the key–query budget across more, narrower heads reduces interference and lowers the required $D_{K}$.
Relation to bounds.
For denser graphs (Figure 12, $r=16$, the empirical thresholds lie slightly below our constructive upper-bound designs—i.e., we need slightly less total key dimension than the construction would guarantee—suggesting either slack in the analysis or other factors the model exploits during training. At the same time, as $r$ grows the gap between the constructive upper bound and the information-theoretic lower bound grows. Together, these trends indicate the true optimum is closer to the lower-bound scaling and that there is room to tighten (or redesign) dense-graph constructions.
Takeaway.
Across densities, capacity remains a threshold phenomenon; the threshold is controlled by the number of edges $m^{\prime}$ rather than the vertices $m$; and the optimal head count scales with $m^{\prime}/d_{\text{model}}$. Practically, when budgeting attention for relational workloads, counting relationships is a more reliable guide than counting vocabulary size. The denser-graph experiments therefore extend the capacity law beyond permutations and provide further evidence for a principled multi-head advantage that grows with graph density.
Figure 11: Sharp transition persists and shifts right for $r\in\{2,4\}$. Mean test micro-F1 vs. total key dimension $D_{K}$ for m=128m{=}128, $d_{\text{model}}{=}32$. More heads reach perfect recovery at smaller $D_{K}$; for $r=2$, a single head plateaus well below 1.0, while for $r=4$ 2 heads has the same property. More heads exhibit a sudden jump to $\approx\!1.0$.
Figure 12: Capacity scales with $m^{\prime}$. Minimal passing $D_{K}$ versus $x=(r\,m/d_{\text{model}})\log_{2}m$ collapses across degrees $r$. The dashed line $D_{K}\approx 0.46\,x$ is a one-parameter fit, highlighting that the number of edges $m^{\prime}$ predicts the threshold more accurately than the vocabulary size $m$. Error bars are calculated using the same methodology as described in Appendix A.2.
Figure 13: Optimal head count rises with density. Number of heads that achieves the smallest passing $D_{K}$ versus graph degree $r$ (here m=128m{=}128, $d_{\text{model}}{=}32$). The trend tracks $m^{\prime}/d_{\text{model}}=(r\,m)/d_{\text{model}}$ up to a constant factor and shows increasing benefit from additional heads as the graph becomes denser. Error bars indicate neighboring head counts that tied for the minimum within the measurement resolution (using the same methodology as described in Appendix A.2).
@@ Appendix B Further Details on Our Explicit Constructions
We here provide additional details on our explicit constructions from Section 4. The computational complexity of all the constructive algorithms we describe in this paper is $O(md_{model}\log m)$. This follows fairly directly from the constructions themselves, although a small bit of care is needed to multiply the matrices together in the right order. This is significantly lower than the computation required to train an attention mechanism using gradient descent. We start with the easiest case - permutation graphs with no embedding. This result is subsummed by the second construction, so is only included as a warm up for the more general case.
@@ B.1 Construction I: Permutation Graphs with One‑Hot Embeddings
3:Construct Key Matrix: Draw $W_{K}\in\mathbb{R}^{m\times d_{k}}$ with i.i.d. entries $(W_{K})_{jl}\sim\text{Bernoulli}(p)$.
4:   For each node $v_{j}$, the key is $\mathbf{k}_{j}=\mathbf{e}_{j}W_{K}$.
5:Construct Query Matrix: For each node $v_{i}$, set its query $\mathbf{q}_{i}=\mathbf{k}_{\pi(i)}$.
6:   This is equivalent to setting the $i$-th row of $W_{Q}$ to be the $\pi(i)$-th row of $W_{K}$.
7:Set Threshold: $\tau=\frac{p+p^{2}}{2}\,d_{k}$.
@@
$\Pr\!\big[S_{i,\pi(i)}\leq\tau\big]\;\leq\;\exp\!\Big(-\tfrac{\mu_{1}\delta_{1}^{2}}{2}\Big)$
where
$\delta_{1}=1-\tfrac{\tau}{\mu_{1}}=\tfrac{1-p}{2},$
@@
$\Pr\!\big[S_{ij}\geq\tau\big]\;\leq\;\exp\!\Big(-\tfrac{\mu_{2}\delta_{2}^{2}}{2+\delta_{2}}\Big)$
where
$\delta_{2}=\tfrac{\tau}{\mu_{2}}-1=\tfrac{1-p}{2p},$
@@
Monotonicity under context restriction immediately now yields correctness for every context, independent of context length. Our lower bound from Section 9 for this case is $\Omega(\frac{m^{\prime}}{d_{\text{model}}}\log m)=\Omega(\log m)$. Our construction achieves an upper bound of $d_{k}=O(\log m)$, demonstrating that the bound is tight for this class of problems. Also note that the threshold proof is identical if softmax is used; see Appendix C.